Circular No. 6789

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971214 R. M. Kippen and P. Woods, University of Alabama in Huntsville

and Marshall Space Flight Center (MSFC); and V. Connaughton, National Academy of Sciences/National Research Council and MSFC, for the CGRO BATSE team; and D. A. Smith, A. M. Levine, and R. Remillard, Massachusetts Institute of Technology, for the RXTE-ASM team; and K. Hurley, University of California at Berkeley, for the Ulysses team, write: “The GRB detected by BeppoSAX (IAUC 6787) also triggered CGRO-BATSE on Dec. 14.97274 UT (trigger no. 6533). The BATSE event consists of a complex series of pulses lasting ∼ 40 s, with an estimated total fluence above 20 keV of 1.09 (± 0.07) × 10^{−5} erg cm^{−2} and a peak flux, occurring at trigger + 13 s, of 1.95 (± 0.05) photons cm^{−2} s^{−1} (50–300 keV; 1.024-s integration). The burst was detected by Ulysses, yielding a preliminary joint BATSE/Ulysses Interplanetary Network (IPN) timing annulus of radius 53^{o}.762, halfwidth 0^{o}.091, and center at α = 11^{h}25^{m}19^{s}, δ = +11^{o}43′.2 (equinox 2000.0). The entire duration of the event was also observed by a single RXTE-ASM camera, which detected a flux enhancement lasting 50 s and reaching a peak intensity of 470 ± 140 mCrab (1.5–12 keV; 1-s integration), 15 s after the BATSE trigger. The singlecamera ASM detection yields a localization region of dimensions 4^{o}.4 × 0^{o}.2, centered at α = 12^{h}03^{m}41^{s}, δ = +64^{o}54′.7, with the following corners: α = 12^{h}20^{m}08^{s}, δ = +63^{o}47′.5; 12^{h}21^{m}04^{s}, +63^{o}57′.5; 11^{h}44^{m}59^{s}, +65^{o}56′.9; 11^{h}45^{m}08^{s}, +65^{o}43′.9. The intersection of the ASM and IPN localizations overlaps the location provided by BeppoSAX given on IAUC 6787. A sky-map of this event showing the various locations is available at

http://www.batse.msfc.nasa.gov/~kippen/batserbr/brbr obs.html.”

XTE J0053–724 D. A. H. Buckley, South African Astronomical Observatory; M. J. Coe

and J. B. Stevens, Southampton University; L. Angelini and N. E. White, Goddard Space Flight Center; and P. Giommi, SAX/SDC, communicate that the SMC source reported on IAUC 6777 and 6788 is in fact the same object as the variable source 1WGA J0053.8–7226, which is identified as the x-ray counterpart to a Be star in the Small Magellanic Cloud. They add: “The optical source has V = 14.9, exhibits strong Hα emission (EW = −1.3 nm), and has a clear infrared excess. Infrared measurements taken in Nov. 1995 and Oct. 1996 show a large infrared increase, probably related to this recent x-ray outburst.”

1997 December 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green